Neutron stars consist only of neutrons and have unbelievably high densities. A typical mass and radius for a neutron star might be 4.1 x 1028 kg and 1.4 x 103 m. (a) Find the density of such a star. (b) If a dime (V = 2.0 x 10-7 m3) were made from this material, how much would it weight (in pounds)?
Answer #1

m = 4.1*10^28 kg, R = 1.4*10^3 m

(a) density = mass/volume

d = 4.1*10^28/((4/3)*3.14*(1.4*10^3)^3)

d = 3.57 *10^18 kg/m^3

(b) V = 2*10^-7 m^3

w = mg = d*V*g

w = 3.57*10^18*2*10^-7*9.8

w = 7*10^12 N

w = 1.574*10^12 lbs